OpenStudy (anonymous):

A street lamp weighs 150 N. It is supported by two wires that form an angle of 120 (degrees) with each other. The tensions in the wires are equal a) what is the tension in each wire? b) if the anggle is reduced to 90 (degrees) what is the tension in each wire? PICTURES WOULD BE REALLY HELPFUL! Thank you :D

OpenStudy (asnaseer):

Part a) is represented by:  \[T\cos(60)+T\cos(60)=150\]